OpenStudy (anonymous):

Find the exat sum or difference. Gerri spends 1/21 of her money on pencils and 2/21 on paper. What fraction of her money does she spend? Give the answer in simplest form. A 1/7 B 3/20 C 4/21 D 8/7

OpenStudy (aripotta):

what's 1/21 + 2/21?

OpenStudy (anonymous):

0.142857143

OpenStudy (aripotta):

or 3/21 lol

OpenStudy (aripotta):

3/21 can be simplified tho. can you tell me what it simplifies to?

sammixboo (sammixboo):

\(\Large \frac{1}{21}+\frac{2}{21}~ \color{red}{->~ }\frac{2+1}{21} ~\color{red}{->~ }\frac{3}{21}\)

OpenStudy (anonymous):

so b

sammixboo (sammixboo):

No :P \(\Large \frac{3}{21}\) can be simplified. We can divide both the numerator and denominator by 3, since 3 is the highest common factor of 3 and 21. \(\Large \frac{3\div 3}{21\div 3}\) Can you tell me what is 3/3 and 21/3?

OpenStudy (fibonaccichick666):

you want to work with fractions here, not a calculator. Make sure you understand how they add together